A pulley has a tangential speed of 14 m/s and an angular velocity of 6.5 rad/sec. What is the normal acceleration of the pulley?
1
Expert's answer
2020-11-17T13:04:05-0500
"r=\frac{v}{\omega}\\a_n=\omega^2r\\a_n=\omega^2\frac{v}{\omega}\\a_n=\omega v=(14)(6.5)=91\frac{m}{s^2}"